# Solving the QCD effective kinetic theory with neural networks

### v1

@@ A.1 Explicit formulas for elastic and inelastic collision kernels
×$\times\Big{(}f(\mathrm{\mathbf{p}})f(\mathrm{\mathbf{k}})(1+f(\mathrm{\mathbf{p^{\prime}}}))(1+f(\mathrm{\mathbf{k^{\prime}}})$ (25)
@@
×$\times\left(\delta^{3}(\mathrm{\mathbf{\tilde{p}}}-\mathrm{\mathbf{p}})+\delta^{3}(\mathrm{\mathbf{\tilde{p}}}-\mathrm{\mathbf{k}})-\delta^{3}(\mathrm{\mathbf{\tilde{p}}}-\mathrm{\mathbf{p^{\prime}}})-\delta^{3}(\mathrm{\mathbf{\tilde{p}}}-\mathrm{\mathbf{k^{\prime}}})\right),$
@@
×$\times\int_{-1}^{1}\mathrm{d}{\cos}\,\theta_{q}\int_{0}^{2\pi}\mathrm{d}{\phi_{q}}\,\int_{0}^{2\pi}\mathrm{d}{\phi}\,_{pq}\int_{0}^{2\pi}\mathrm{d}{\phi}\,_{kq},$
@@
[…]
@@ A.2 Monte Carlo evaluation of the collision terms
$v^{\mu}\partial_{\mu}n_{ijk}$ $=\int\frac{\mathrm{d}^{3}{\mathrm{\mathbf{\tilde{p}}}}\,}{(2\pi)^{3}}w_{ijk}(\mathrm{\mathbf{\tilde{p}}})v^{\mu}\partial_{\mu}f(\mathrm{\mathbf{\tilde{p}}})$ (33)
$=-\int\frac{\mathrm{d}^{3}{\mathrm{\mathbf{\tilde{p}}}}\,}{(2\pi)^{3}}w_{ijk}(\tilde{\mathrm{\mathbf{p}}})\mathcal{C}[f(\mathrm{\mathbf{\tilde{p}}})]$ (34)
@@ Appendix B Restframe
is diagonal. In the rest frame, it can be decomposed in its eigenvectors $\{u_{\nu}^{\mu}\}$ (lower index labeling the eigenvectors), where its diagonal components give the energy density $\varepsilon$ and the pressures $P_{i}$ in different directions. Importantly, $u_{0}$ is the only timeline eigenvector, which is normalized to $u_{0}^{2}=-1$, while the others are spacelike $u_{i}^{2}=1$. Note that we use the mostly-plus metric convention here, $\eta_{\mu\nu}=\mathrm{diag}(-1,1,1,1)$. Diagonalizing a matrix is a well-known problem in linear algebra. In our case, the eigenvector equation is given by
@@
However, this Lorentz transformation is not unique. This is because an eigenvector is only defined up to a multiplicative constant, and even a normalized eigenvector can still be multiplied by a factor $(-1)$. We will now discuss how to fix the transformation uniquely. First, we require that $P^{\mu}{}_{\nu}$ is an orthochronous Lorentz transformation, which means that it does not flip time (time reversal). The condition for that is that $P^{0}{}_{0}>0$. If $P^{0}{}_{0}<0$, we multiply the timelike eigenvector with a factor $(-1)$, or, equivalently, perform
@@
After that, one might additionally perform a parity transformation to enforce positive determinant, $\det P\neq+1$. For that, one could multiply all spacelike eigenvectors with $(-1)$, or, equivalently, perform the parity transformation
@@
is positive for all spatial directions. This can be achieved by fixing the signs in (44), or, equivalently, multiplying every basis vector by $(-1)$ that corresponds to a direction of negative particle current. Clearly, the current (45) transforms Lorentz covariantly, because of the Lorentz invariant phase space measure $\mathrm{d}^{3}{\mathrm{\mathbf{p}}}\,/p_{0}$ and the Lorentz vector $p^{\mu}$. As is evident from (38), for upper indices we need to use $P^{-1}$. Thus, for any Lorentz transformation $P$, we can easily obtain $J^{\mu}$ in the rest frame.

### v2

@@ A.1 Explicit formulas for elastic and inelastic collision kernels
$\times\Big(f(\mathrm{\mathbf{p}})f(\mathrm{\mathbf{k}})(1+f(\mathrm{\mathbf{p^{\prime}}}))(1+f(\mathrm{\mathbf{k^{\prime}}})$ (24)
@@
$\times\left(\delta^{3}(\mathrm{\mathbf{\tilde{p}}}-\mathrm{\mathbf{p}})+\delta^{3}(\mathrm{\mathbf{\tilde{p}}}-\mathrm{\mathbf{k}})-\delta^{3}(\mathrm{\mathbf{\tilde{p}}}-\mathrm{\mathbf{p^{\prime}}})-\delta^{3}(\mathrm{\mathbf{\tilde{p}}}-\mathrm{\mathbf{k^{\prime}}})\right),$
@@
$\times\int_{-1}^{1}\mathrm{d}{\cos}\,\theta_{q}\int_{0}^{2\pi}\mathrm{d}{\phi_{q}}\,\int_{0}^{2\pi}\mathrm{d}{\phi}\,_{pq}\int_{0}^{2\pi}\mathrm{d}{\phi}\,_{kq},$
@@
$\begin{split}\mathcal{C}^{1\leftrightarrow 2}[f(\mathrm{\mathbf{\tilde{p}}})]&=\frac{(2\pi)^{3}}{4\pi\tilde{p}^{2}}\frac{1}{\nu}\int_{0}^{\infty}\mathrm{d}{p}\,\int_{0}^{p/2}\mathrm{d}{k^{\prime}}\,4\pi\gamma^{p}_{p^{\prime},k^{\prime}}\\ &\quad\times\Big\{f(\mathrm{\mathbf{p}})(1+f(p^{\prime}\hat{\mathrm{\mathbf{p}}}))(1+f(k^{\prime}\hat{\mathrm{\mathbf{p}}}))\\ &\qquad\quad-f(p^{\prime}\hat{\mathrm{\mathbf{p}}})f(k^{\prime}\hat{\mathrm{\mathbf{p}}})(1+f(\mathrm{\mathbf{p}}))\Big\}\\ &\times\left[\delta(\tilde{p}-p)-\delta(\tilde{p}-p^{\prime})-\delta(\tilde{p}-k^{\prime})\right]\end{split}$ (28)
with $p^{\prime}=p-k^{\prime}$ and $\hat{\mathrm{\mathbf{p}}}=\mathrm{\mathbf{p}}/p$ being a unit vector, which enforces that all particles are collinear in the direction of $\mathrm{\mathbf{p}}$. The effective splitting/joining rates $\gamma^{p}_{p^{\prime}k}$ interpolate between the Bethe-Heitler and LPM regime and are obtained via Arnold et al. (2003a)
@@ A.2 Monte Carlo evaluation of the collision terms
[…]
@@ Appendix B Restframe
is diagonal. It can be decomposed in its eigenvectors $\{u_{\nu}^{\mu}\}$ (lower index labeling the eigenvectors), where its diagonal components give the energy density $\varepsilon$ and the pressures $P_{i}$ in different directions. Importantly, $u_{0}$ is the only timeline eigenvector, which is normalized to $u_{0}^{2}=-1$, while the others are spacelike $u_{i}^{2}=1$. Note that we use the mostly-plus metric convention here, $\eta_{\mu\nu}=\mathrm{diag}(-1,1,1,1)$. Diagonalizing a matrix is a well-known problem in linear algebra. In our case, the eigenvector equation is given by
@@
However, this Lorentz transformation is not unique. This is because an eigenvector is only defined up to a multiplicative constant, and even a normalized eigenvector can still be multiplied by a factor (1)(-1). We will now discuss how to fix the transformation uniquely. First, we require that $P^{\mu}{}_{\nu}$ is an orthochronous Lorentz transformation, which means that it does not flip time (time reversal). The condition for that is that $P^{0}{}_{0}>0$. If $P^{0}{}_{0}<0$, we multiply the timelike eigenvector with a factor (1)(-1), or, equivalently, perform
@@
After that, one might additionally perform a parity transformation to enforce positive determinant, $\det P\neq+1$. For that, one could multiply all spacelike eigenvectors with (1)(-1), or, equivalently, perform the parity transformation
@@
is positive for all spatial directions. This can be achieved by fixing the signs in (43), or, equivalently, multiplying every basis vector by (1)(-1) that corresponds to a direction of negative particle current. Clearly, the current (44) transforms Lorentz covariantly, because of the Lorentz invariant phase space measure $\mathrm{d}^{3}{\mathrm{\mathbf{p}}}\,/p_{0}$ and the Lorentz vector $p^{\mu}$. As is evident from (37), for upper indices we need to use $P^{-1}$. Thus, for any Lorentz transformation $P$, we can easily obtain $J^{\mu}$ in the rest frame.